# Echoes from the Dark: Galaxy Catalog Incompleteness in Standard Siren Cosmology

### v1

@@ 2.4 Implementation in CHIMERA
The completeness function is computed in multiple $N_{\text{masks}}$ masks following the approach of Finke et al. (2021). Healpix pixels are grouped together using a k-means algorithm based on the number of galaxies inside each pixel. For each mask, the completeness function (Eq. 11) is evaluated by binning the observed galaxies into redshift bins and comparing their density to a theoretical background distribution. As for the catalog term, the completeness function can be saved to disk and reloaded to be later interpolated onto the specific redshift grid where the likelihood is evaluated. Note that the pixelization scheme used for the completeness can be coarser than that of the GW data. In this work, we adopt the default settings of CHIMERA ($N_{\text{masks}}=5$, nside = 32, and 100 redshift bins) that have been chosen to ensure adequate statistics within each mask.
@@
The background galaxy distribution is obtained by integrating the Schechter mass function over the relevant host mass range with any applied host weighting (Eq. 12). This term is used in both the (pre)computation of the completeness function $P_{\rm comp}$ and, crucially, is evaluated at each likelihood call and therefore must be defined efficiently, as it directly impacts the overall computational performance of the pipeline. In CHIMERA, this is optimized using just-in-time (JIT) compilation and vectorization across all redshift grids. An example illustrating the background galaxy distributions is shown in Figure 1c.
@@ 3 Catalogs
This section presents the generation of mock galaxy and GW catalogs. In particular, we design these catalogs to explore the precision achievable in $H_{0}$ at various incompleteness levels for different weighting schemes, assuming the host properties are well known. To facilitate comparison with the previous analysis in B24, we adopt the same mock galaxy catalog and assumptions for the GW population. Further details on the framework are provided in Appendix A.
@@ 3.1 Parent galaxy catalog
We select a subsample of MICECATv2 so that the redshift distribution can be described with a parametric function and efficiently computed at each likelihood call of CHIMERA. We first compute the stellar mass distribution of the original catalog, defining the minimum mass as the mass corresponding to the peak density at each redshift bin (Fig. 2, left panel). We then model the galaxy stellar mass distribution using a double Schechter function that evolves with redshift (Fig. 2, central panel), where the redshift evolution is driven by the evolution of the minimum mass. Lastly, we subsample the original catalog so that our model can describe the background galaxy distribution of the catalog with a typical percentage difference below $5\,\%$ (Fig. 2, right panel). The complete procedure is detailed in Appendix A. This allows us to flexibly describe the redshift distribution of our parent sample and easily extend it with weighting while keeping a realistic clustering of galaxies. The resulting catalog, identified as the parent galaxy catalog throughout this paper, contains approximately 335 million galaxies (about $67\%$ of MICECATv2) and is more than 200 times larger than the high-mass subsample of MICECATv2 galaxies used in B24.
@@ 3.2 Potential host galaxies
In this study, we assign the probability that each galaxy hosts a BBH merger as a function of its stellar mass. We consider three weighting schemes:
$w\propto M_{\star}^{\alpha_{M}}\quad\text{with}\quad\alpha_{M}=\begin{cases}0&(\texttt{mock0})\\ 1&(\texttt{mock1})\\ 2&(\texttt{mock2})\end{cases}$ (16)
These cases are compared in Fig. 3, alongside the equivalent weighting adopted in B24 and the low-$z$ trend from the population synthesis study from Artale et al. (2020). Despite the uniform ($\alpha_{M}=0$) weighting being maximally agnostic, it unrealistically assigns equal hosting probability to all galaxies, regardless of their stellar mass, across a range spanning several orders of magnitude. Applying a stellar-mass cut is a practical attempt to mitigate this issue. The $\alpha_{M}=1$ weighting is more in line with theoretical predictions in the local universe, where most of the BBH events are observed (see Sect. 3.4). Finally, the $\alpha_{M}=2$ weighting allows us to explore the case of an even stronger dependence on stellar mass.
We remark that the parent catalog is complete above a given minimum stellar mass that evolves with redshift (Fig. 2). This allows us to test the realistic scenario where the density of galaxies is highest and therefore less informative for the standard siren method. This occurs for mock0 at low $z$, where the events simulated in this work that are more informative for $H_{0}$ are present. However, caution is required when using mock0 to constrain $H(z)$ to higher $z$ or, more importantly, when using $\alpha_{M}<0$, as the galaxy density of this parent sample may not necessarily reflect a realistic density. This issue does not affect mocks generated with $\alpha_{M}>0$, since the weights preferentially select high-mass galaxies where the completeness of the parent sample is highest.
Figure 3: Host galaxy weight as a function of stellar mass for different weighting schemes: constant weighting at $\log_{10}(M_{\star}/M_{\odot})>10.5$ (solid line, as in Borghi et al. 2024), linear (dashed), and quadratic (dotted) weighting. The gray line illustrates the $z=0.1$ trend from Artale et al. (2020).
@@ 3.3 Mock GW catalogs
We populate our three host galaxy catalogs with BBH mergers now assigned independently of individual galaxy properties. We assume that the overall BBH merger rate follows the cosmic star formation rate density from Madau & Dickinson (2014):
@@
where $\psi(z)\propto(1+z)^{\gamma}$ at low $z$, peaks at $z_{\rm p}$, and then declines as $\psi(z)\propto(1+z)^{-\kappa}$. The redshifts are then converted to luminosity distances by assuming the cosmology of MICECATv2. For the mass distribution, we adopt the phenomenological model “PowerLaw+Peak” (PLP, Talbot & Thrane 2018), which is the preferred model up to LVK GWTC-3 (Abbott et al. 2023b). Equation (4) is then factorized as follows:
@@
Figure 4: Main properties of 100 O5-like BBH events with $\mathrm{S/N}>25$, drawn from MICECATv2 with three host weighting schemes: unweighted (mock0, gray), weighted by stellar mass (mock1, blue), and weighted by stellar mass squared (mock2, red). Top panels: Distribution in the sky localization ($90\%$ C.L.) versus luminosity-distance uncertainty plane color-coded by the network $\mathrm{S/N}$. Bottom panels: Distributions of source redshift, number of galaxies within the localization volume ($N_{\mathrm{gal,vol}}$), host galaxy stellar mass, and $i$-band magnitude for the various samples; the arrows indicate the median values. The dot-dashed lines show the background redshift distribution from which GW events are drawn. The vertical dashed line shows the mass and magnitude cuts that are applied in this work to study incompleteness effects.
@@ 3.4 Detected GW events
To simulate detections, we use GWFAST (Iacovelli et al. 2022a, b). We assume quasi-circular non-precessing BBH systems, so that the following detector frame parameters characterize each waveform:
@@
We select a subsample of 100 events with a network signal-to-noise ratio, $\mathrm{S/N}>25$. These cuts are designed to yield the 100 best events for each configuration over approximately one year of observation, assuming host galaxies are uniformly distributed in comoving volume, consistently with B24. The properties of the resulting catalogs are summarized in Fig. 4. The top panels show the sky localization area at $90\,\%$ C.L. ($\Delta\Omega_{90\%}$) versus luminosity distance error ($\Delta d_{L}$), color-coded by $\mathrm{S/N}$. The bottom panels show redshift distributions, galaxy counts within the localization volume ($N_{\mathrm{gal,vol}}$, computed assuming a flat prior on $H_{0}\in[60,80]\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$), and host galaxy properties: stellar mass $M_{\star}$ and $i$-band magnitude.
The observed GW events are located at $z\lesssim 1$, with sky localization areas ranging from $0.1$ to a few tens of square degrees. The values of $N_{\mathrm{gal,vol}}$, a key indicator of how strongly each GW event can constrain $H_{0}$ in the dark siren regime, range from $\mathcal{O}(10^{2})$ to $\mathcal{O}(10^{5})$, with the maximum set by the finite redshift coverage of the current catalog. These values are significantly larger than those in B24, where the best-localized events contained only a few galaxies and could therefore be treated as “pseudo-bright” sirens. Differences among the three mocks reflect both the background host galaxy distribution (see Fig. 4, bottom left panel) and the specific realization of the sampled GW events. Due to the lower redshift distribution of their hosts, mock0 events are observed with slightly higher $\mathrm{S/N}$ and better sky localization, with median values of $34.1$ and $1.3\penalty 10000\ \mathrm{deg^{2}}$, respectively, compared to mock1 ($32.5$ and $1.8\penalty 10000\ \mathrm{deg^{2}}$) and mock2 ($31.6$ and $2.0\penalty 10000\ \mathrm{deg^{2}}$). As a result, the median $N_{\mathrm{gal,vol}}$ differs between the mocks: $\approx 6\,500$ for mock0 and $\approx 12\,000$ for both mock1 and mock2.
@@ 3.5 Observed galaxy catalogs
Figure 5: Effects of galaxy survey incompleteness on $H_{0}$ constraints for different stellar mass (left panel) and $i$-band magnitude (right panel) completeness cuts across three host weighting schemes: unweighted (mock0, grey), weighted by stellar mass (mock1, blue), and weighted by stellar mass squared (mock2, red).
@@ 4.1 Complete galaxy catalog
When the galaxy catalog contains all potential hosts, we obtain percent-level constraints on $H_{0}$. The precision improves when host weighting favors more massive galaxies, with $\sigma_{H_{0}}=1.6\%,\,1.3\%,0.9\%$ for mock0, mock1, mock2, respectively. The main reason for this improvement is the decrease in the effective number of galaxies within the GW localization volume $N_{\rm gal,vol}$ that significantly contribute to the catalog term once the weighting is applied. We verify that the trend is independent of the specific realization of GW events by performing one-dimensional analyses on the $H_{0}$ posterior, keeping all the other hyperparameters fixed to their fiducial values. This is further supported by the fact that mock0, despite having lower $N_{\rm gal,vol}$ values (see Fig. 4, bottom panels), provides the weakest constraints.
These results are remarkable given that typical $N_{\rm gal,vol}$ values are $200\times$ larger than in B24. Even with this increased galaxy density, constraints degrade by only a factor of two in the least informative case We attribute this to the fact that, even when the analysis is extended to more realistic and denser catalogs, the redshift uncertainties provided by a spectroscopic galaxy catalog remain sufficiently small to preserve the information content of the catalog term (Eq. 7). Consequently, through the combined analysis of multiple events, the true value of $H_{0}$ can still be statistically recovered with high precision.
@@ 4.3 Modeling incompleteness effects on H0
To describe how galaxy catalog incompleteness affects the precision of $H_{0}$ constraints, we model the evolution of the percentage uncertainty $\sigma_{H_{0}}$ using a sigmoid function that smoothly connects the case of a complete catalog ($\sigma_{H_{0}}^{\rm comp}$) to the empty catalog/spectral siren limit ($\sigma_{H_{0}}^{\rm spec}$):
@@
where $x$ is the chosen completeness parameter (such as stellar mass or magnitude), $x_{\mathrm{th}}$ is the threshold at which the uncertainty most rapidly transitions between the two regimes, and $k$ sets the steepness of the transition.
For mass-completeness cuts, we find $x_{\rm th}\equiv\log_{10}M_{\rm th}=10.8,\ 11.2,\ 11.6$ and k=4.2, 11.1, 22.4k=4.2,\ 11.1,\ 22.4 for mock0, mock1, and mock2, respectively. As $\alpha_{M}$ increases, the threshold shifts toward higher stellar masses, indicating that missing less massive (less probable) hosts ultimately has a negligible impact on the $H_{0}$ constraints. Notably, with mass and mass-squared weighting, percent-level precision is still achievable as long as the catalog is complete at $\log_{10}(M_{\star}/M_{\odot})>11$. The annotations in Fig. 5 show the number of true host galaxies remaining after each completeness cut. Percent-level constraints are obtained when at least about $20$ events have their true hosts inside the catalog. Interestingly, there is one configuration (mock2 with $\log_{10}(M_{\rm cut}/M_{\odot})=11.5$) for which this is possible thanks to a single lucky event. We underscore the importance of spectroscopic follow-ups to target the localization region of these best-constraining “golden” events with future galaxy surveys.
@@
$\langle P_{\rm compl}\rangle\equiv\frac{1}{V_{\rm c}(z_{\rm hor})}\int_{0}^{z_{\rm hor}}P_{\rm compl}(z)\,\mathrm{d}V_{\rm c}\quad\text{with}\quad z_{\rm hor}=1.4$ (23)

### v2

@@ 2.4 Implementation in CHIMERA
The completeness function is computed in multiple $N_{\text{masks}}$ masks following the approach of Finke et al. (2021). Healpix pixels are grouped together using a k-means algorithm based on the number of galaxies inside each pixel. For each mask, the completeness function (Eq. 11) is evaluated by binning the observed galaxies into redshift bins and comparing their density to a theoretical background distribution. As for the catalog term, the completeness function can be saved to disk and reloaded to be later interpolated onto the specific redshift grid where the likelihood is evaluated. Note that the pixelization scheme used for the completeness can be coarser than that of the GW data. In this work, we adopt the default settings ($N_{\text{masks}}=5$, nside = 32, and 100 redshift bins) that have been chosen to ensure adequate statistics within each mask.
@@
The background galaxy distribution is obtained by integrating the Schechter mass function over the relevant host mass range with any applied host weighting (Eq. 12). This term is used in both the (pre)computation of the completeness function, $P_{\rm comp}$, and, crucially, is evaluated at each likelihood call and therefore must be defined efficiently, as it directly impacts the overall computational performance of the pipeline. In CHIMERA, this is optimized using just-in-time (JIT) compilation and vectorization across all redshift grids. An example illustrating the background galaxy distributions is shown in Figure 1c.
@@ 3 Catalogs
This section describes the generation of mock galaxy and GW catalogs. First, we build our parent galaxy catalog, which represents all galaxies in our synthetic universe and is described by a parametric, redshift-evolving stellar mass function (Sect. 3.1). From this parent catalog, we derive three sub-catalogs of potential BBH hosts by applying different host weighting schemes (Sect. 3.2). For each host catalog, we generate mock GW events following a prescribed population model (Sect. 3.3) and simulate their detection with a Fisher matrix approach (Sect. 3.4). Finally, we apply observational completeness cuts to the parent galaxy catalog to simulate realistic survey conditions (Sect. 3.5). The parent galaxy catalog, either complete or with cuts applied, serves as the input catalog for the parameter estimation runs. Further details on this framework are provided in Appendix A.
@@ 3.1 Parent galaxy catalog
We select a subsample of MICECATv2 so that the redshift distribution can be described with a parametric function and efficiently computed at each likelihood call of CHIMERA, while preserving realistic galaxy clustering. We first compute the stellar mass distribution of the original catalog, defining the minimum mass as the mass corresponding to the peak density at each redshift bin (Fig. 2, left panel). We then model this distribution with a redshift-evolving double Schechter function (Fig. 2, central panel), where the evolution is driven by the minimum mass. Lastly, we subsample the original catalog so that our model can describe the background galaxy distribution of the catalog with a typical percentage difference below $5\,\%$ (Fig. 2, right panel). To compute the theoretical number density (Eq. 12) the Schechter function is integrated from a redshift-evolving minimum mass to $\log_{10}(M_{\star}/M_{\odot})\approx 12$. Further details are given in Appendix A. The resulting catalog, identified as the parent galaxy catalog throughout this paper, contains approximately 335 million galaxies (about $67\%$ of MICECATv2) and is more than 200 times larger than the high-mass subsample of MICECATv2 galaxies used in B24.
@@ 3.2 Potential host galaxies
In this study, we assign the probability that each galaxy hosts a BBH merger as a function of its stellar mass without applying mass cuts. We consider three weighting schemes:
$w\propto M_{\star}^{\alpha_{M}},\quad\text{where}\quad\alpha_{M}=0,\,1,\,2.$ (16)
These cases are compared in Fig. 3 alongside the weighting adopted in B24 and the low-$z$ trend from the population synthesis study of Artale et al. (2020). Despite the uniform ($\alpha_{M}=0$) weighting being maximally agnostic, it unrealistically assigns equal hosting probability to all galaxies, regardless of their stellar mass, across a range spanning several orders of magnitude. Applying a stellar-mass cut is a practical attempt to mitigate this issue. The $\alpha_{M}=1$ weighting is more in line with theoretical predictions in the local Universe, where most of the BBH events are observed (see Sect. 3.4). Finally, the $\alpha_{M}=2$ weighting allows us to explore the case of an even stronger dependence on stellar mass.
We remark that the parent catalog is complete above a given minimum stellar mass that evolves with redshift (Fig. 2). In the likelihood, we integrate the Schechter function over this mass range to ensure statistical consistency. This allows us to probe the regime in which the galaxy density is highest and therefore least informative for the dark siren method. In practice, this occurs for $\alpha_{M}=0$ at low $z$, where most of the events are located. However, since low-mass galaxies are increasingly missed at higher redshift, caution would be required if extending this analysis to constrain $H(z)$ at higher $z$ with the $\alpha_{M}=0$ catalog (or even more so with $\alpha_{M}<0$), as the galaxy density of this parent sample may not necessarily reflect the true underlying population. This issue does not affect mocks generated with $\alpha_{M}>0$, since the weights preferentially select high-mass galaxies where the fidelity of the parent sample is highest.
Figure 3: Host galaxy weighting schemes considered in this work: constant (gray), linear (blue), and quadratic (red) mass weighting. For comparison, we show the $z=0.1$ trend from Artale et al. (2020) (dotted blue line) and the cut used in previous work (solid line, as in Borghi et al. 2024).
@@ 3.3 Mock GW catalogs
For each weighting scheme, we populate the corresponding host galaxy catalog with BBH mergers, generating three mock GW catalogs hereafter labeled as mock0, mock1, mock2 for $\alpha_{M}=0,1,2$ respectively. We assume that the overall BBH merger rate follows the cosmic star formation rate density from Madau & Dickinson (2014):
@@
where $\psi(z)\propto(1+z)^{\gamma}$ at low $z$, peaks at $z_{\rm p}$, and then declines as $\psi(z)\propto(1+z)^{-\kappa}$. The redshifts are then converted to luminosity distances by assuming the cosmology of MICECATv2. For the mass distribution, we adopt the phenomenological model “PowerLaw+Peak” (PLP, Talbot & Thrane 2018), which is the preferred model up to LVK GWTC-3 (Abbott et al. 2023b). Equation (4) is then factorized as
@@
Figure 4: Main properties of 100 O5-like BBH events with $\mathrm{S/N}>25$ drawn from MICECATv2 with three host weighting schemes: unweighted (mock0, gray), weighted by stellar mass (mock1, blue), and weighted by stellar mass squared (mock2, red). Top panels: Distribution in the sky localization ($90\%$ C.L.) versus luminosity-distance uncertainty plane color-coded by the network $\mathrm{S/N}$. Bottom panels: Distributions of source redshift, number of galaxies within the localization volume ($N_{\mathrm{gal,vol}}$), host galaxy stellar mass, and $i$-band magnitude for the various samples; the arrows indicate the median values. The dot-dashed lines show the background redshift distribution of potential BBH host galaxies used to generate GW events. The vertical dashed line shows the mass and magnitude cuts that are applied in this work to study incompleteness effects.
@@ 3.4 Detected GW events
To simulate detections, we use GWFAST (Iacovelli et al. 2022a, b). We assume quasi-circular non-precessing BBH systems so that the following detector frame parameters characterize each waveform:
@@
We select a subsample of 100 BBHs with a network signal-to-noise ratio, $\mathrm{S/N}>25$. This cut, chosen to be consistent with B24, is designed to yield the 100 best BBH events over approximately one year of observation. In general, higher S/N events have tighter sky localization and fewer potential host galaxies, resulting in stronger $H_{0}$ constraints in the dark siren regime. In contrast, pure spectral siren analyses (empty catalog limit), included here for comparison, would benefit from including lower S/N events to better resolve features in the mass distribution (see e.g., Abbott et al. 2023b).
Figure 4 presents the main properties of the 100 BBH events generated for each of the three host weighting schemes. The top panels show the $90\,\%$ C.L. sky localization area against the luminosity distance uncertainty, color-coded by $\mathrm{S/N}$. The bottom panels show the redshift distribution, galaxy counts within the localization volume ($N_{\mathrm{gal,vol}}$), and host galaxy properties (stellar mass $M_{\star}$ and $i$-band magnitude). Note that $N_{\mathrm{gal,vol}}$ assumes a flat prior on $H_{0}\in[60,80]\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$. The observed GW events are located at $z\lesssim 1$, with sky localization areas ranging from $0.1$ to a few tens of square degrees. The values of $N_{\mathrm{gal,vol}}$, a key indicator of how strongly each GW event can constrain $H_{0}$ in the dark siren regime, range from $\mathcal{O}(10^{2})$ to $\mathcal{O}(10^{5})$, with the maximum set by the finite redshift coverage of the current catalog. These values are significantly larger than those in B24, where the best-localized events contained only a few galaxies and could therefore be treated as “pseudo-bright” sirens. Differences among the three mocks reflect both the background host galaxy distribution (see Fig. 4, bottom left panel) and the specific realization of the sampled GW events. Due to the lower redshift distribution of their hosts, mock0 events are observed with slightly higher $\mathrm{S/N}$ and better sky localization, with median values of $34.1$ and $1.3\penalty 10000\ \mathrm{deg^{2}}$, respectively, compared to mock1 ($32.5$ and $1.8\penalty 10000\ \mathrm{deg^{2}}$) and mock2 ($31.6$ and $2.0\penalty 10000\ \mathrm{deg^{2}}$). As a result, the median $N_{\mathrm{gal,vol}}$ differs between the mocks: $\approx 6\,500$ for mock0 and $\approx 12\,000$ for both mock1 and mock2.
@@ 3.5 Observed galaxy catalogs
Figure 5: Effects of galaxy survey incompleteness on $H_{0}$ constraints for different stellar mass (left panel) and $i$-band magnitude (right panel) completeness cuts across three host weighting schemes: unweighted (mock0, gray), weighted by stellar mass (mock1, blue), and weighted by stellar mass squared (mock2, red). The values annotated below each panel indicate the number of host galaxies retained in the catalog after applying each cut.
@@ 4.1 Complete galaxy catalog
When the galaxy catalog contains all potential hosts, we obtain percent-level constraints on $H_{0}$. The precision improves when host weighting favors more massive galaxies, with $\sigma_{H_{0}}=1.6\%,\,1.3\%,0.9\%$ for mock0, mock1, mock2, respectively. The main reason for this improvement is the decrease in the effective number of galaxies within the GW localization volume, $N_{\rm gal,vol}$, that significantly contribute to the catalog term once the weighting is applied. We verify that the trend is independent of the specific realization of GW events by performing one-dimensional analyses on the $H_{0}$ posterior, keeping all the other hyperparameters fixed to their fiducial values. This is further supported by the fact that mock0, despite having lower $N_{\rm gal,vol}$ values (see Fig. 4, bottom panels), provides the weakest constraints.
These results are remarkable given that typical $N_{\rm gal,vol}$ values are $200\times$ larger than in B24. Even with this increased galaxy density, constraints degrade by only a factor of two in the least informative case. We attribute this to the fact that, even when the analysis is extended to more realistic and denser catalogs, the redshift uncertainties provided by a spectroscopic galaxy catalog remain sufficiently small to preserve the information content of the catalog term (Eq. 7). Consequently, through the combined analysis of multiple events, the true value of $H_{0}$ can still be statistically recovered with high precision.
@@ 4.3 Modeling incompleteness effects on H0
To describe how galaxy catalog incompleteness affects the precision of $H_{0}$ constraints, we model the evolution of the percentage uncertainty, $\sigma_{H_{0}}$, using a sigmoid function that smoothly connects the case of a complete catalog ($\sigma_{H_{0}}^{\rm comp}$) to the empty catalog/spectral siren limit ($\sigma_{H_{0}}^{\rm spec}$):
@@
where $x$ is the chosen completeness parameter (such as stellar mass or magnitude), $x_{\mathrm{th}}$ is the threshold at which the uncertainty transitions most rapidly between the two regimes, and $k$ sets the steepness of the transition.
For mass-completeness cuts, we find $x_{\rm th}\equiv\log_{10}M_{\rm th}=10.8,\ 11.2,\ 11.6$ and k=4.2, 11.1, 22.4k=4.2,\ 11.1,\ 22.4 for mock0, mock1, and mock2, respectively. As $\alpha_{M}$ increases, the threshold shifts toward higher stellar masses, indicating that missing less massive (less probable) hosts ultimately has a negligible impact on the $H_{0}$ constraints. Notably, with mass and mass-squared weighting, percent-level precision is still achievable as long as the catalog is complete at $\log_{10}(M_{\star}/M_{\odot})>11$. The annotations in Fig. 5 show the number of true host galaxies remaining after each completeness cut. Percent-level constraints are obtained when at least about $20$ events retain their hosts after completeness cuts or, equivalently, when the catalog completeness is high enough to keep a sufficient fraction of true host galaxies in the catalog. Interestingly, there is one configuration (mock2 with $\log_{10}(M_{\rm cut}/M_{\odot})=11.5$) for which this is possible thanks to a single lucky event. We underscore the importance of spectroscopic follow-ups to target the localization region of these best-constraining “golden” events with future galaxy surveys.
@@
$\langle P_{\rm compl}\rangle\equiv\frac{1}{V_{\rm c}(z_{\rm hor})}\int_{0}^{z_{\rm hor}}P_{\rm compl}(z)\,\mathrm{d}V_{\rm c}\quad\text{with}\quad z_{\rm hor}=1.4\,$, (23)